Find the slope of the line that passes through the points (2,7) and (4,2)
Answer by mananth(16946)

x1 y1 x2 y2
2 7 4 2

slope m = (y2-y1)/(x2-x1)
( 2 - 7 )/( 4 - 2 )
( -5 / 2 )
m= -5/2
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 2 , 7 ) in
Y = m x + b
7.00 = -5/2 * 2 + b
7.00 = -5 + b
b= 7 + 5
b= 12
So the equation will be
Y = -5/2 x + 12
m.ananth@hotmail.ca